a father is 25 years older than his son. The sum of their ages is 53 years. Find their ages

Answer by Ash14(5)

let age of father=x
and age of his son =y
now
father is 25 years older than his son.
means that
x-y=25........(1)
and
the sum of their ages is 53 years
means
x+y=53.......(2)
now from (1) and (2)
x+y=53
x-y=25
============ (cancel y because y-y=0 and x+x=2x)
2x=78
x=78/2
x=39
sub value of x in (1)
x-y=25
(39)-y=25(here we need to make y the subject)
|
|
39=25+y
39-25=y
14=y
so
x=39(age of father) and y=14(age of son)
.......
*******
.......
good luck :)